no. of absences, x 0 1 2 3 4 5 6 7 8 9
Final exam score, y 88.3 85.5 82.8 81.2 78.3 73.4 64.4 70.9 64.7 66.4
n
3 .997
4 .950
5 .878
6 .811
7 .754
8 .707
9 .666
10 .632
11 .602
12 .576
13 .553
14 .532
15 .514
16 .497
17 .482
18 .468
19 .456
20 .444
21 .433
22 .423
23 .413
24 .404
25 .396
26 .388
27 .381
28 .374
29 .367
30 .361

The accompanying data represent the number of days​ absent, x, and the final exam​ score, y, for a sample of college students in a general education course at a large state university. Complete parts ​(a) through​ (e) below.

a) y = __x + __

b) interpret the slope and the y-intercept, if appropriate

c) predict the final exam score for a student who missed four and five class periods
Answer #1

Ans:

a)

Regression equation:

y'=-2.7727 x+88.067

b)slope indicate that for each absence there will be decrease of on average 2.7727 scores in final exam.

y-intercept indicates that if there is no absence,average score is 88.067

c)

For x=4,y'=-2.7727*4+88.067=76.98

For x=5,y'=-2.7727*5+88.067=74.20